if I contain 3 moles of gas in a container with a volume of 60 liters and at a temperature of 400K, what is the pressure inside the container?



1
Expert's answer
2022-05-12T00:39:02-0400

PV=nRT

P=nRT/V

n=3 mole

V=60 l=0.06 m3

R=8.31 J/(mole K)

T=400 K

P=3(8.31)400/0.06=166200 Pa